Consider a population with a mean of 70 and a standard deviation of 6. A random sample of size
36 is drawn. What is the probability that the sample mean is between 70:5 and 71:5 ?
1
Expert's answer
2020-03-30T12:05:04-0400

If "n>30," the Central Limit Theorem can be used.

Let "X_1,X_2,...,X_n" be a random sample from a distribution with mean "\mu" and variance "\sigma^2." Then if "n" is sufficiently large, has approximately a normal distribution with "\mu_{\bar{X}}=\mu" and "\sigma_{\bar{X}}=\sigma^2/n."

Give that "n=36, \mu=70, \sigma=6."

"n=36>30=>X\sim(\mu;\sigma^2/n)". Then


"Z={\bar{X}-\mu \over \sigma/\sqrt{n}}\sim N(0;1)"

"P(70.5<\bar{X}<71.5)=P(Z<{71.5-70\over 6/\sqrt{36}})-P(Z<{70.5-70\over 6/\sqrt{36}})="

"=P(Z<{1.5})-P(Z<0.5)\approx0.93319-0.69146\approx0.2417"



The probability that the sample mean is between "70.5" and "71.5" is "0.2417."